# DreamMotion: Space-Time Self-Similarity Score Distillation for Zero-Shot Video Editing

### v1

@@ Abstract
Text-driven diffusion-based video editing presents a unique challenge not encountered in image editing literature: establishing real-world motion. Unlike existing video editing approaches, here we focus on score distillation sampling to circumvent the standard reverse diffusion process and initiate optimization from videos that already exhibit natural motion. Our analysis reveals that while video score distillation can effectively introduce new content indicated by target text, it can also cause significant structure and motion deviation. To counteract this, we propose to match space-time self-similarities of the original video and the edited video during the score distillation. Thanks to the use of score distillation, our approach is model-agnostic, which can be applied for both cascaded and non-cascaded video diffusion frameworks. Through extensive comparisons with leading methods, our approach demonstrates its superiority in altering appearances while accurately preserving the original structure and motion. Project page: https://hyeonho99.github.io/dreammotion/
@@ 1 Introduction
Building upon the progress in diffusion models [46, 14, 49], the advent of large-scale text-image pairs [42] brought unprecedented breakthrough in text-driven image generative tasks. In particular, real-world image editing has undergone significant evolution, supported by foundational Text-to-Image (T2I) diffusion models [40, 39, 41, 34]. However, extending the success of diffusion-based image editing to video editing introduces a significant challenge: modelling temporally consistent, real-world motion throughout the reverse diffusion process.
@@
Existing methods leveraging T2I diffusion models typically start by inflating attention layers to attend to multiple frames simultaneously [60, 22, 56, 37, 7, 61, 65, 8, 21]. Yet, this technique falls short in achieving smooth and complete motion, as it depends on the implicit preservation of motion through the inflated attention layers. As a result, commonly adopted solution is to employ additional visual hints that explicitly guide the reverse diffusion process. One strategy is to use attention map guidance, for example, by injecting self-attention maps [4, 37] or manipulating cross-attentions [28]. Other works attempt to integrate the denoising process with spatially-aligned structural cues, like depth or edge maps. For example, pretrained adapter networks such as ControlNet [64] or GLIGEN [26] have been transferred from image to video domain, achieving structure-consistent outputs [7, 65, 17, 21].
@@
Even with the presence of pretrained Text-to-Video (T2V) diffusion models, zero-shot video editing still poses a significant hurdle since publicly available T2V models [55, 50] lack a sufficiently rich temporal prior to accurately depict real-world motion in the generated videos, as illustrated in Fig. 2. Thus, recent endeavors often adopt a self-supervised strategy of finetuning pretrained model weights on the motion presented in an input video [68, 20, 59, 66]. Whether employing T2I or T2V models, the conventional reverse diffusion process —beginning with standard Gaussian noise or, at most, inverted latent representations— struggles to reprogram complex, real-world motion, unless supplemented by additional visual conditions or by overfitting the spatial-temporal prior to a particular video.
@@
To this end, we propose to diverge from the previous video editing literature. Our approach, DreamMotion, deliberately avoids the standard denoising process (ancestral sampling), and instead leverages the Score Distillation Sampling (SDS, [36]) grounded optimization to edit a video. Specifically, starting from an input video with temporally-consistent, natural motion, we attempt to progressively modify the video’s appearance while maintaining the integrity of the motion. In specific, our framework gradually injects target appearance to the video using Delta Denoising Score (DDS, [11]) gradients within T2V diffusion models. During this procedure, we filter the gradients with additional binary mask conditions to avoid blurriness and over-saturation. While this optimization effectively infuses the targeted appearance, it tends to accumulate structural errors, resulting in deviations in motion across the final output frames (see Fig. 3). To address this, we present self-similarity-based space-time regularization methods. More specifically, by aligning the spatial self-similarity of diffusion features between the original and edited videos, we preserve structure and motion integrity while seamlessly modifying the appearance. Furthermore, ensuring temporal self-similarity between the two features facilitates effective temporal smoothing, preventing potential distortions in areas subjected to optimization. Our methodology is applied to both cascaded and non-cascaded video diffusion models, showcasing its wide applicability across different video editing frameworks.
In summary, our work offers the following key contributions:
@@
  • A pioneering zero-shot framework that distills video score from text-to-video diffusion prior to inject target appearance.
@@
Figure 2: Ancestral sampling based zero-shot video editing fails to capture complex, real-world motion in the generated videos.
@@ 2.0.2 Conditional Generation
In the context of conditional generation, data $\boldsymbol{x}$ is paired with an additional conditioning signal $y$, which in our case is a text caption. To train a text-driven diffusion model, the text conditional embedding $y$ is incorporated to the objective as:
@@ 2.0.3 Video Diffusion Models
Our framework leverages foundational video diffusion models for obtaining video score. Consider a video sequence of $N$ frames represented by $\boldsymbol{x}^{1:N}\in\mathbb{R}^{N\times d}$. For any $n$-th frame within this sequence, denoted by $\boldsymbol{x}^{n}\in\mathbb{R}^{d}$, the noisy frame latent $\boldsymbol{x}_{t}^{n}$ sampled from $p_{t}(\boldsymbol{x}_{t}^{n}|\boldsymbol{x}^{n})$ can be expressed as $\boldsymbol{x}_{t}^{n}=\sqrt{\bar{\alpha}_{t}}\boldsymbol{x}^{n}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}_{t}^{n}$ , where $\boldsymbol{\epsilon}_{t}^{n}\sim\mathcal{N}(0,I)$. Then, we similarly define $\boldsymbol{x}_{t}^{1:N}$ and $\boldsymbol{\epsilon}_{t}^{1:N}$. The objective of video diffusion model training is then to obtain a denoiser network $\epsilon_{\phi^{*}}$ that satisfies:
@@
Seeking to create videos that are both spatially and temporally enlarged and of high quality, video diffusion models have been expanded to cascaded pipelines [16, 13, 45, 2, 57, 63]. These cascaded video pipelines commonly follow a coarse-to-fine video generation approach, beginning with a module dedicated to creating keyframes that are low in both spatial and temporal resolution. Subsequent stages involve temporal interpolation and spatial super-resolution modules, which work to increase the temporal and spatial resolution of the frames, respectively. In this work, we plug our method to both cascaded and non-cascaded scenarios, proving its model-agnostic capability.
@@ 3.1 Overview
Starting with a series of input video frames $\boldsymbol{\hat{x}}^{1:N}$, a corresponding text prompt $\hat{y}$, and a target text $y$, our goal is to get an edited video $\boldsymbol{x}^{1:N}$ that preserves the structural integrity and overall motion of $\boldsymbol{\hat{x}}^{1:N}$, while faithfully reflecting $y$. DreamMotion starts by initializing the target video variable $\boldsymbol{x}_{0}^{1:N}(\theta)$ by the original video $\boldsymbol{\hat{x}}^{1:N}$. Our optimization strategy is then three-pronged: (1) $\mathcal{L}_{\text{V-DDS}}$ that paints $\boldsymbol{x}_{0}^{1:N}(\theta)$ to match the appearance dictated by $y$, (2) $\mathcal{L}_{\text{S-SSM}}$ which encourages the structure of $\boldsymbol{x}_{0}^{1:N}(\theta)$ to align with $\boldsymbol{\hat{x}}^{1:N}$, (3) $\mathcal{L}_{\text{T-SSM}}$ which smoothens the gradients over temporal dimension to eliminate any potential artifacts.
In Sec. 3.2, we briefly review SDS and DDS loss formulations, and describe how we directly modify appearance of $\boldsymbol{x}^{1:N}$ with DDS-based gradients. This technique, while effective in appearance injection, tends to accumulate structural inaccuracies, resulting in motion deviation in the end output. To address this, Sec. 3.3 introduces a strategy for structural correction based on self-similarity, and Sec. 3.4 details our approach for temporal smoothing, also leveraging self-similarity. Finally, in Sec. 3.5, we elaborate on the extension of DreamMotion to the cascaded video diffusion framework. For simplicity, we primarily describe the diffusion model as operating in pixel space throughout this paper. However, in practice, our implementation encompasses both a latent-space based (Sec. 4.1, [50]) and a pixel-space based video diffusion model (Sec. 4.2, [63]).
@@ 3.2.1 Image Score Distillation
Under the assumption that the SDS score should be zero for pairs of correctly matched prompts and images, DDS [11] enhances the gradient direction obtained from the SDS framework by incorporating an additional text-image pair, comprising a reference text $\boldsymbol{\hat{y}}$ and a reference image $\boldsymbol{\hat{x}}_{0}$. Specifically, the noisy direction of SDS score is calculated using the reference text-image branch, and this noisy score is subtracted from the main SDS optimization branch:
@@ 3.2.2 Video Score Distillation with Masked Gradients
Leveraging a pretrained T2V diffusion model $\boldsymbol{\epsilon}_{\phi}$, we extend the DDS mechanism to distill video scores. Let $\boldsymbol{x}_{0}^{1:N}(\theta)$ represent the target video parameterized by $\theta$, and $\boldsymbol{x}_{0}^{1:N}$ denote the fixed, source video. We optimize the video variable $\boldsymbol{x}_{0}^{1:N}(\theta)$ to reflect target text $y$ by minimizing:
@@ 3.3.1 Spatial Self-Similarity Matching
where $cos(\cdot,\cdot)$ denotes the normalized cosine similarity and $\boldsymbol{x}_{t}^{1:N}(\theta)$ is simplified to $\boldsymbol{x}_{t}^{1:N}$ for brevity. The spatial self-similarity matching objective is formulated as:
@@
Figure 6: Filtering optimization gradients plays a crucial role in maintaining visual fidelity and preserving structure of the input video. See Fig. 12 for full results and masks.
@@ 3.4.1 Temporal Self-Similarity Matching
Calculating self-similarity over time necessitates a method to compress spatial information while retaining essential spatial details. For this purpose, we employ spatial marginal mean a first-order statistic, spatial marginal mean, as our global descriptor. This choice is supported by prior works [24, 62], which have demonstrated its effectiveness in capturing crucial spatial details and serving as a robust global descriptor. More concretely, we condense the spatial dimensions of the extracted key features $K(\boldsymbol{x}_{t}^{1:N})\in\mathbb{R}^{N\times(H\times W)\times C}$ to $M[K(\boldsymbol{x}_{t}^{1:N})]\in\mathbb{R}^{N\times C}$ through the process defined as:
@@
$TS_{i,j}(\boldsymbol{x}_{t}^{1:N})=cos(M_{i}[K(\boldsymbol{x}_{t}^{1:N})],M_{j}[K(\boldsymbol{x}_{t}^{1:N})]).$ (14)
Subsequently, the temporal self-similarity matching loss is formulated as:
@@
Figure 7: Comparison. DreamMotion, applied to Zeroscope model, is evaluated against five baseline methods. For a detailed assessment, visit our project page.
@@ 4.1.1 Setup
For evaluation, we choose 26 text-video pairs from the public DAVIS [35] and WebVid [1] datasets. The videos vary in length from 8 frames to 16 frames. In this experiment, we deploy our method on ZeroScope [50], a foundational text-to-video latent diffusion model. The CFG scale $w$ is configured as 9.0. We perform optimization for 200 steps using stochastic gradient descent (SGD) with learning rate of 0.4. The optimization of an 8-frame video requires approximately 2 minutes, while optimizing a 16-frame video takes around 4 minutes, utilizing a single A100 GPU.
@@ 4.1.4 Quantitative Results
We conducted a comprehensive quantitative evaluation, which includes both automatic metrics and a user study. The summarized results can be found in Tab. 1. (a) Automatic metrics. We employ CLIP [38] for automatic evaluation. To assess textual alignment [12], we measure the average cosine similarity between the textual prompt and the frames that have been edited. In terms of frame consistency, we calculate CLIP image features for every frame in the output video and then compute the average cosine similarity across all pairs of frames within the video. According to the results in Tab. 1, our approach surpasses the baselines in achieving higher textual alignment and better temporal consistency. (b) User study. We surveyed 36 participants to assess the accuracy of editing, temporal consistency, and preservation of structure & motion, using a rating scale from 1 to 5. Participants were shown the input video followed by anonymized output videos from each baseline. They were then asked the three questions: (i) Edit Accuracy: Does the output video accurately reflect the target text by appropriately editing all relevant elements? (ii) Frame Consistency: Are the frames in the output video temporally consistent? (iii) Structure and Motion Preservation: Has the structure and motion of the input video been accurately maintained in the output video? Tab. 1 illustrates that our method outperforms the baselines in all measured aspects.
Automatic Metric Human Eval.
@@
TAV 0.8177 0.9218 3.52 2.82 2.89
CV 0.7850 0.9678 2.74 2.68 2.03
CAV 0.7848 0.9297 2.17 2.16 2.18
Gen-1 0.8192 0.9704 3.31 3.62 2.95
Tokenflow 0.7813 0.9576 3.63 3.54 3.92
Ours (Zeroscope) 0.8209 0.9726 4.14 4.21 4.33
Inv + Word Swap 0.7586 0.9714 3.36 3.42 2.21
@@
Table 1: Quantitative evaluations. Our method, when applied on Zeroscope, outperforms non-cascaded baselines, while our approach applied on Show-1 surpasses cascaded baselines, in all five features.
Figure 8: Comparison. Our approach, applied to Show-1 cascaded model, is evaluated against two baselines.

### v2

# Project page: https://hyeonho99.github.io/dreammotion

@@ Abstract
Text-driven diffusion-based video editing presents a unique challenge not encountered in image editing literature: establishing real-world motion. Unlike existing video editing approaches, here we focus on score distillation sampling to circumvent the standard reverse diffusion process and initiate optimization from videos that already exhibit natural motion. Our analysis reveals that while video score distillation can effectively introduce new content indicated by target text, it can also cause significant structure and motion deviation. To counteract this, we propose to match the space-time self-similarities of the original video and the edited video during the score distillation. Thanks to the use of score distillation, our approach is model-agnostic, which can be applied for both cascaded and non-cascaded video diffusion frameworks. Through extensive comparisons with leading methods, our approach demonstrates its superiority in altering appearances while accurately preserving the original structure and motion.
Keywords:
Video Editing Diffusion Models Score Distillation
@@ 1 Introduction
Building upon the progress in diffusion models [49, 14, 52], the advent of large-scale text-image pairs [45] brought an unprecedented breakthrough in text-driven image generative tasks. In particular, real-world image editing has undergone significant evolution, supported by foundational Text-to-Image (T2I) diffusion models [43, 42, 44, 37]. However, extending the success of diffusion-based image editing to video editing introduces a significant challenge: modeling temporally consistent, real-world motion throughout the reverse diffusion process.
@@
Existing methods leveraging T2I diffusion models typically start by inflating attention layers to attend to multiple frames simultaneously [63, 22, 59, 40, 7, 64, 68, 8, 21]. Yet, this technique falls short of achieving smooth and complete motion, as it depends on the implicit preservation of motion through the inflated attention layers. As a result, a commonly adopted solution is to employ additional visual hints that explicitly guide the reverse diffusion process. One strategy is to use attention map guidance, for example, by injecting self-attention maps [4, 40] or manipulating cross-attentions [29]. Other works attempt to integrate the denoising process with spatially-aligned structural cues, like depth or edge maps. For example, pre-trained adapter networks such as ControlNet [67] or GLIGEN [27] have been transferred from image to video domain, achieving structure-consistent outputs [7, 68, 17, 21].
@@
Even with the presence of pre-trained Text-to-Video (T2V) diffusion models, zero-shot video editing still poses a significant hurdle since publicly available T2V models [58, 53] lack sufficiently rich temporal priors to accurately depict real-world motion in the generated videos, as illustrated in Fig. 2. Thus, recent endeavors often adopt a self-supervised strategy of finetuning pre-trained model weights on the motion presented in an input video [71, 20, 35, 62, 69]. Whether employing T2I or T2V models, the conventional reverse diffusion process —beginning with standard Gaussian noise or, at most, inverted latent representations— struggles to reprogram complex, real-world motion, unless supplemented by additional visual conditions or by overfitting the spatial-temporal priors to a particular video.
@@
To this end, we propose to diverge from the previous video editing literature. Our approach, DreamMotion, deliberately avoids the standard denoising process (ancestral sampling), and instead leverages the Score Distillation Sampling (SDS, [39]) grounded optimization to edit a video. Specifically, starting from an input video with temporally consistent, natural motion, we attempt to progressively modify the video’s appearance while maintaining the integrity of the motion. In specific, our framework gradually injects target appearance to the video using Delta Denoising Score (DDS, [11]) gradients within T2V diffusion models. During this procedure, we filter the gradients with additional binary mask conditions to avoid blurriness and over-saturation. While this optimization effectively infuses the targeted appearance, it tends to accumulate structural errors, resulting in deviations in motion across the final output frames (see Fig. 3). To address this, we present self-similarity-based space-time regularization methods. More specifically, by aligning the spatial self-similarity of diffusion features between the original and edited videos, we preserve structure and motion integrity while seamlessly modifying the appearance. Furthermore, ensuring temporal self-similarity between the two features facilitates effective temporal smoothing, preventing potential distortions in areas subjected to optimization. Our methodology is applied to both cascaded and non-cascaded video diffusion models, showcasing its wide applicability across different video editing frameworks.
In summary, DreamMotion offers the following key contributions:
@@
  • A pioneering zero-shot framework that distills video score from text-to-video diffusion priors to inject target appearance.
@@
Figure 2: Ancestral sampling-based zero-shot video editing fails to capture complex, real-world motion in the generated videos.
@@ 2.0.2 Conditional Generation
In the context of conditional generation, data $\boldsymbol{x}$ is paired with an additional conditioning signal $y$, which in our case is a text caption. To train a text-driven diffusion model, the text conditional embedding $y$ is incorporated into the objective as:
@@ 2.0.3 Video Diffusion Models
Our framework leverages foundational video diffusion models for obtaining video scores. Consider a video sequence of $N$ frames represented by $\boldsymbol{x}^{1:N}\in\mathbb{R}^{N\times d}$. For any $n$-th frame within this sequence, denoted by $\boldsymbol{x}^{n}\in\mathbb{R}^{d}$, the noisy frame latent $\boldsymbol{x}_{t}^{n}$ sampled from $p_{t}(\boldsymbol{x}_{t}^{n}|\boldsymbol{x}^{n})$ can be expressed as $\boldsymbol{x}_{t}^{n}=\sqrt{\bar{\alpha}_{t}}\boldsymbol{x}^{n}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}_{t}^{n}$ , where $\boldsymbol{\epsilon}_{t}^{n}\sim\mathcal{N}(0,I)$. Then, we similarly define $\boldsymbol{x}_{t}^{1:N}$ and $\boldsymbol{\epsilon}_{t}^{1:N}$. The objective of video diffusion model training is then to obtain a denoiser network $\epsilon_{\phi^{*}}$ that satisfies:
@@
Seeking to create videos that are both spatially and temporally enlarged and of high quality, video diffusion models have been expanded to cascaded pipelines [16, 13, 48, 2, 60, 66]. These cascaded video pipelines commonly follow a coarse-to-fine video generation approach, beginning with a module dedicated to creating keyframes that are low in both spatial and temporal resolution. Subsequent stages involve temporal interpolation and spatial super-resolution modules, which work to increase the temporal and spatial resolution of the frames, respectively. In this work, we plug our method into both cascaded and non-cascaded scenarios, proving its model-agnostic capability.
@@ 3.1 Overview
Starting with a series of input video frames $\boldsymbol{\hat{x}}^{1:N}$, a corresponding text prompt $\hat{y}$, and a target text $y$, our goal is to get an edited video $\boldsymbol{x}^{1:N}$ that preserves the structural integrity and overall motion of $\boldsymbol{\hat{x}}^{1:N}$, while faithfully reflecting $y$. DreamMotion starts by initializing the target video variable $\boldsymbol{x}_{0}^{1:N}(\theta)$ by the original video $\boldsymbol{\hat{x}}^{1:N}$. Our optimization strategy is then three-pronged: (1) $\mathcal{L}_{\text{V-DDS}}$ that paints $\boldsymbol{x}_{0}^{1:N}(\theta)$ to match the appearance dictated by $y$, (2) $\mathcal{L}_{\text{S-SSM}}$ which encourages the structure of $\boldsymbol{x}_{0}^{1:N}(\theta)$ to align with $\boldsymbol{\hat{x}}^{1:N}$, (3) $\mathcal{L}_{\text{T-SSM}}$ which smoothens the gradients over the temporal dimension to eliminate any potential artifacts.
In Sec. 3.2, we briefly review SDS and DDS loss formulations and describe how we directly modify the appearance of $\boldsymbol{x}^{1:N}$ with DDS-based gradients. This technique, while effective in appearance injection, tends to accumulate structural inaccuracies, resulting in motion deviation in the end output. To address this, Sec. 3.3 introduces a strategy for structural correction based on self-similarity, and Sec. 3.4 details our approach for temporal smoothing, also leveraging self-similarity. Finally, in Sec. 3.5, we elaborate on the extension of DreamMotion to the cascaded video diffusion framework. For simplicity, we primarily describe the diffusion model as operating in pixel space throughout this paper. However, in practice, our implementation encompasses both a latent space-based (Sec. 4.1, [53]) and a pixel space-based video diffusion model (Sec. 4.2, [66]).
@@ 3.2.1 Image Score Distillation
Under the assumption that the SDS score should be zero for pairs of correctly matched prompts and images, DDS [11] enhances the gradient direction obtained from the SDS framework by incorporating an additional text-image pair, comprising a reference text $\boldsymbol{\hat{y}}$ and a reference image $\boldsymbol{\hat{x}}_{0}$. Specifically, the noisy direction of the SDS score is calculated using the reference text-image branch, and this noisy score is subtracted from the main SDS optimization branch:
@@ 3.2.2 Video Score Distillation with Masked Gradients
Leveraging a pre-trained T2V diffusion model $\boldsymbol{\epsilon}_{\phi}$, we extend the DDS mechanism to distill video scores. Let $\boldsymbol{x}_{0}^{1:N}(\theta)$ represent the target video parameterized by $\theta$, and $\boldsymbol{x}_{0}^{1:N}$ denote the fixed, source video. We optimize the video variable $\boldsymbol{x}_{0}^{1:N}(\theta)$ to reflect target text $y$ by minimizing:
@@ 3.3.1 Spatial Self-Similarity Matching
where $cos(\cdot,\cdot)$ denotes the normalized cosine similarity, $i,j$ are all pairs of spatial indexes $(1\leq i,j\leq(H{\times}W))$, and $\boldsymbol{x}_{t}^{1:N}(\theta)$ is simplified to $\boldsymbol{x}_{t}^{1:N}$ for brevity. The spatial self-similarity matching objective is formulated as:
@@
Figure 6: Filtering optimization gradients plays a crucial role in maintaining visual fidelity and preserving the structure of the input video. Bounding boxes detected by off-the-shelf models [26, 30] are used to create binary masks indicating the target regions for editing.
@@ 3.4.1 Temporal Self-Similarity Matching
Calculating self-similarity over time necessitates a method to compress spatial information while retaining essential spatial details. For this purpose, we employ spatial marginal mean a first-order statistic, spatial marginal mean, as our global descriptor. This choice is supported by prior works [24, 65], which have demonstrated their effectiveness in capturing crucial spatial details and serving as a robust global descriptor. More concretely, we condense the spatial dimensions of the extracted key features $K(\boldsymbol{x}_{t}^{1:N})\in\mathbb{R}^{N\times(H\times W)\times C}$ to $M[K(\boldsymbol{x}_{t}^{1:N})]\in\mathbb{R}^{N\times C}$ through the process defined as:
@@
$TS_{i,j}(\boldsymbol{x}_{t}^{1:N})=cos(M_{i}[K(\boldsymbol{x}_{t}^{1:N})],M_{j}[K(\boldsymbol{x}_{t}^{1:N})]),$ (14)
where $i,j$ are from frame indexes $(1\leq i,j\leq N)$. Subsequently, the temporal self-similarity matching loss is formulated as:
@@
Figure 7: Comparison. DreamMotion, applied to the Zeroscope model, is evaluated against five baseline methods. For a detailed assessment, please visit our project page.
@@ 4.1.1 Setup
For evaluation, we chose 26 text-video pairs from the public DAVIS [38] and WebVid [1] datasets. The videos vary in length from 8 frames to 16 frames. In this experiment, we deploy our method on ZeroScope [53], a foundational text-to-video latent diffusion model. The CFG scale $w$ is configured as 9.0. We perform optimization for 200 steps using stochastic gradient descent (SGD) with a learning rate of 0.4. The optimization of an 8-frame video requires approximately 2 minutes, while optimizing a 16-frame video takes around 4 minutes, utilizing a single A100 GPU.
@@ 4.1.4 Quantitative Results
We conducted a comprehensive quantitative evaluation, which includes both automatic metrics and a user study. The summarized results can be found in Tab. 1.
(a) Automatic metrics. We first employ CLIP [41] to measure the text alignment and frame consistency of the edited videos. For assessing textual alignment [12], we measure average cosine similarity between the target text prompt and the edited frames. In terms of frame consistency, we calculate CLIP image features for every frame in the output video and then compute the average cosine similarity across all neighboring pairs of frames. We additionally compute tracking-based motion fidelity score [65] and framewise LPIPS [29] for measuring spatial consistency. According to the results in Tab. 1, our approach surpasses the baselines in achieving higher textual alignment and better spatial-temporal consistency.
(b) User study. We surveyed 36 participants to assess the accuracy of editing, temporal consistency, and preservation of structure & motion, using a rating scale from 1 to 5. Participants were shown the input video followed by anonymized output videos from each baseline. They were then asked the three questions: (i) Edit Accuracy: Does the output video accurately reflect the target text by appropriately editing all relevant elements? (ii) Frame Consistency: Are the frames in the output video temporally consistent? (iii) Structure and Motion Preservation: Has the structure and motion of the input video been accurately maintained in the output video? Tab. 1 illustrates that our method outperforms the baselines in all measured aspects.
Automatic Metrics Human Evaluation
Method Text-Align Frame-Con Motion-Fidelity Frame-LPIPS Edit-Acc Frame-Con SM-Preserve
Tune-A-Video 0.8177 0.9218 0.6947 0.4172 3.52 2.82 2.89
ControlVideo 0.7850 0.9678 - 0.3763 2.74 2.68 2.03
Control-A-Video 0.7848 0.9297 0.8453 0.3829 2.17 2.16 2.18
Gen-1 0.8192 0.9704 - - 3.31 3.62 2.95
Tokenflow 0.7813 0.9576 0.9184 0.3427 3.63 3.54 3.92
Ours (Zeroscope) 0.8209 0.9726 0.9259 0.3042 4.14 4.21 4.33
Table 1: Quantitative evaluations. DreamMotion with Zeroscope outperforms various video editing methods in all seven features.
Figure 8: Comparison. DreamMotion with Show-1 cascaded model is evaluated against two baselines.
Automatic Metrics Human Evaluation
@@
Inversion + Word Swap 0.7586 0.9714 3.36 3.42 2.21
@@
Table 2: Quantitative evaluations. DreamMotion utilizing Show-1 surpasses other cascaded baselines across the five features. Other baselines were also implemented using the same video model, ensuring a fair comparison.